OpenStudy (vera_ewing):

When figures (including points) are rotated 270° counterclockwise about the origin, it is also the same rotating figures clockwise by what other degree amount? @wio

OpenStudy (vera_ewing):

90 degrees right?

OpenStudy (anonymous):

 In general, you do \(360 - \theta = \theta'\)

OpenStudy (vera_ewing):

oh so the answer is 360 degrees NOT 90 degrees?

OpenStudy (anonymous):

no, see, \(360 - 270 = \ldots\)

OpenStudy (vera_ewing):

Point (5, -2) is reflected across the line y = x. Where is the new location?

OpenStudy (anonymous):

That should transform it \((y,x)\)

OpenStudy (anonymous):

one is asking about a point, the answers are talking about triangles. that makes no sense.

OpenStudy (vera_ewing):

oops sorry lol i don't have answer choices :O

OpenStudy (vera_ewing):

@wio what's the answer? is it (2,5) ???

OpenStudy (anonymous):

no

OpenStudy (vera_ewing):

"That should transform it (y,x)" <---- what does that mean? is that the answer? @wio

OpenStudy (vera_ewing):

@wio ?

OpenStudy (anonymous):

it means that \((x,y)\) becomes \((y,x)\).

OpenStudy (vera_ewing):

ok so answer is (-2, 5) ?

OpenStudy (anonymous):

yes

OpenStudy (vera_ewing):

thanks